### v1

# MDCure: A Scalable Pipeline for Multi-Document Instruction-Following

Abstract
Multi-document (MD) processing is crucial for LLMs to handle real-world tasks such as summarization and question-answering across large sets of documents. While LLMs have improved at processing long inputs, MD contexts still present challenges, such as managing inter-document dependencies, redundancy, and incoherent structures. We introduce MDCure, a scalable and effective fine-tuning pipeline to enhance the MD capabilities of LLMs without the computational cost of pre-training or reliance on human annotated data. MDCure is based on generation of high-quality synthetic MD instruction data from sets of related articles via targeted prompts. We further introduce MDCureRM, a multi-objective reward model which filters generated data based on their training utility for MD settings. With MDCure, we fine-tune a variety of LLMs, from the FlanT5, Qwen2, and LLAMA3.1 model families, up to 70B parameters in size. Extensive evaluations on a wide range of MD and long-context benchmarks spanning various tasks show MDCure consistently improves performance over pre-trained baselines and over corresponding base models by up to 75.5%. Our code, datasets, and models are available at https://github.com/yale-nlp/MDCure.
MDCure: A Scalable Pipeline for Multi-Document Instruction-Following
Gabrielle Kaili-May Liu1 Bowen Shi1 Avi Caciularu2 Idan Szpektor2  Arman Cohan1,3 1Yale University  2Google Research  3Allen Institute for AI kaili.liu@yale.edu     arman.cohan@yale.edu
1 Introduction
With the rapid expansion of bodies of online texts in domains of science, finance, education, news, and more, multi-document (MD) processing becomes a critical aspect of LLMs’ ability to understand, leverage, and reason over large quantities of text information, with applications such as (query-focused) multi-document summarization (MDS) (Fabbri et al., 2019; Xiao et al., 2022; Giorgi et al., 2023), multi-hop question-answering (MHQA) (Yang et al., 2018; Tu et al., 2019), and cross-document coreference resolution (Cattan et al., 2021b; Eirew et al., 2022).
While LLMs continue to upscale context size and improve performance on complex tasks, their ability to perform multi-document tasks is still limited (Hosseini et al., 2024; Lior et al., 2024; Wang et al., 2024d). MD processing presents unique challenges such as managing redundancy, high inter-document relatedness, contradictions, omissions, and lack of coherent inter-document structure (Radev, 2000; Hendrickx et al., 2009; Hewapathirana et al., 2023), which are exacerbated when documents are long (Liu et al., 2024a). While there is substantial work on increasing the long-context processing abilities of LLMs (e.g., An et al. (2024b); Gao et al. (2024a); Xiao et al. (2024)), increased context (Lior et al., 2024; Tang et al., 2024) and model size (Wang et al., 2024d) often does not have substantial impacts on MD capabilities.
Figure 1: The MDCure pipeline generates diverse multi-document instructions, filters them using fine-grained scoring from MDCureRM, and tunes a base LLM to enhance its multi-document capabilities.
Continued pre-training of long-context LMs on sets of related documents has proven effective for MD processing (Caciularu et al., 2021; Yasunaga et al., 2022; Caciularu et al., 2023). However, such methods are not extensible to broader tasks and often require large amounts of pre-training data (e.g., Caciularu et al., 2021; Xiao et al., 2022; Caciularu et al., 2023; Peper et al., 2024).
We propose MDCure, a scalable synthetic instruction-following data generation approach to adapt LLMs for MD settings. Inspired by recent work showing the value of small-scale, high-quality supervised data for efficiently improving LMs’ downstream task performance (Zhou et al., 2023; Sachdeva et al., 2024; Tirumala et al., 2024), MDCure addresses the challenges of data scarcity, complexity, and long-context processing in MD tasks. Instead of relying on costly (Cobbe et al., 2021; Yue et al., 2024) and potentially inadequate human-annotated instructions (Srivastava et al., 2023; Chen et al., 2024c) we leverage powerful LLMs such as GPT-4 (OpenAI et al., 2024) to generate and filter synthetic instructions.
MDCure automatically curates high-quality MD instruction data to enhance LLMs’ MD capabilities. It divides the process into two phases: Generation, where zero-shot prompt templates are used to create complex, cross-text instructions from a set of related documents provided as input context, and Filtering, where a multi-objective, MD-specific reward model, MDCureRM, assesses the generated instructions based on key criteria. This filtering ensures high-quality, diverse, and MD-focused instructions. We find that MDCureRM is pivotal to improving the constitution of the resulting instruction dataset. Instead of generating documents synthetically, by grounding the instructions in real-world texts, MDCure produces realistic and relevant instruction data, improving the model’s ability to synthesize information across multiple sources.
We showcase the efficacy of MDCure through extensive experiments on 6 multi-document benchmarks in over 6 content domains. To analyze the impact of MDCure at scale, we create MDCure instruction datasets of size 12K, 36K, and 72K to fine-tune FlanT5 (Chung et al., 2022), Qwen2-Instruct (Yang et al., 2024), and LLAMA3.1-Instruct (Dubey et al., 2024) models up to 70B parameters in size. Our results show that despite its simplicity, MDCure enables models to consistently achieve superior performance versus existing MD pre-training approaches and outperform prior instruction-following LLMs in both zero- and few-shot settings, with gains of up to 75.5% average improvement. Our contributions are as follows:
  • MDCure, a scalable and general procedure for creating high-quality multi-document instructions that target cross-document relationships and are grounded in real-world contexts, which is robust across model sizes and types.
  • A suite of instruction data complementary to collections such as FLAN (Longpre et al., 2023) for improving MD task performance.
  • A novel set of criteria by which to effectively assess the utility of instruction samples for improving MD performance.
  • MDCureRM, a reward model used as an evaluator specifically designed for the MD setting, to filter and select high-quality MD instruction data in a cost-effective way.
2 Related Work
Multi-Document Modeling. A common approach for MD modeling with LLMs is to perform flat concatenation of all input documents to process with a long-context model (e.g., Longformer (Beltagy et al., 2020), LongT5 (Guo et al., 2022), CoLT5 (Ainslie et al., 2023)). However, such models lack sufficient skill at synthesizing cross-document information (Caciularu et al., 2022; Wolhandler et al., 2022; Chen et al., 2023; Hosseini et al., 2024). More recently, pre-training-based approaches have proven valuable and involve (continual) training of long-context LMs on instances formulated over sets of related documents. Methods such as token unmasking and masked sentence generation have been used to enhance cross-document information utilization for summarization and classification tasks (Zhang et al., 2020; Caciularu et al., 2021; Yasunaga et al., 2022; Xiao et al., 2022). QAMDen (Caciularu et al., 2023) extends these works with a pre-training objective focused on cross-document question-answering. PELMS (Peper et al., 2024) subsequently improves MDS performance by using semantic clustering to determine salience and intentional sentence removal to boost abstractiveness. Other notable recent works leverage LLMs to better perform targeted content selection for MDS and QMDS (Kurisinkel and chen, 2023; Kurisinkel and Chen, 2023; Li et al., 2023a). In contrast to such methods, our approach employs synthetic instruction generation to overcome limitations of heuristic-based data generation and the computational expense of pre-training, offering a scalable way to adapt LLMs for a wide variety of MD tasks.
Synthetic Instruction Generation. Instruction tuning (IT) is a commonly used approach to align pre-trained LMs for various tasks by training on supervised instruction-answer pairs (Sanh et al., 2022; Wei et al., 2022). However, the scarcity of domain-specific instruction data remains a challenge, as manual annotations are costly and impractical at scale (Song et al., 2024; Ziegler et al., 2024). To address this, recent works have focused on leveraging LLMs for synthetic data generation (Honovich et al., 2023; Wang et al., 2023; Chen et al., 2024a). Most of these efforts target general single-document tasks, but some, such as Chen et al. (2024c), delve into MHQA by generating and merging single-hop QA pairs. While effective, these approaches often focus on MHQA and rely on complex frameworks. Similarly, Lupidi et al. (2024) generates synthetic MHQA data but is limited to two-hop tasks. In contrast, MDCure targets broader MD tasks, with varying complexity, providing a simpler, more scalable solution. Data filtering is also critical, as synthetic data can be noisy or nonfactual (Yue et al., 2024; Ziegler et al., 2024). Several works employ LLMs for selecting quality data at scale (Du et al., 2023; Li et al., 2024), with fine-grained rewards used to enhance task-specific filtering (Wang et al., 2024c). Inspired by these approaches, we design MDCureRM, a novel multi-objective MD-specific reward model to score the diversity and quality of MD instruction data, which we use to refine initially generated instruction pools.
3 MDCure
MDCure (Fig. 1) utilizes a two-step approach to produce high-quality MD instruction samples over sets of related documents. These samples can be used via instruction tuning to improve base LLMs’ MD modeling capabilities.
3.1 Generation Phase
In the Generation phase, we sample from a set of carefully tailored zero-shot templates (App. A.2) to prompt a black-box LLM to generate candidate instructions for a given set of documents. The aim is to produce relevant instruction-answer pairs concerning information spanning multiple documents in the context of the LLM. Versus Caciularu et al. (2023), our instructions are more diverse and open-ended. These prompt templates are designed to encourage cross-document reasoning by requiring answers that synthesize information from multiple documents or text segments. They encourage the use of multiple sources instead of only one document, thus involving deeper connections among the provided texts. The wide variety of templates ensures diverse task formulations, ranging from yes/no answers to detailed summaries, which together help generate synthetic MD data that reflects real-world complexity and encourages LLMs to develop strong cross-document comprehension. We analyze the effect of different prompt approaches on the quality of resulting generations in §5.4.
Dataset Source Selection. MDCure’s generation focuses on constructing candidate instructions from sets of documents, which may be sourced from existing datasets of related texts or created by clustering general text corpora. In our experiments, we use topically related document sets from the news domain to model cross-text relationships, building on prior work that shows training with related document sequences can enhance model performance (Caciularu et al., 2021, 2023; Shi et al., 2023). Nonetheless, our generation procedure is equally functional over sets of unrelated documents, which may be utilized to target creation of instruction samples concerning processing of conflicting or irrelevant cross-document information. We base our generated instructions on the NewSHead dataset (Gu et al., 2020). Additional details are in App. A.1.
Generator Model. MDCure is not generally dependent on any specific generator LLM. In combination with our proposed filtering procedure (§3.2), any strong instruction-following LLM can be used to construct effective instruction data. We utilize GPT-3.5-Turbo in our experiments to balance generation quality and costs. [^1] Other cost efficient LLMs such as GPT-4o-mini or Gemini 1.5 Flash can also be used; our generation pipeline was done before the release of these more cost-efficient models. Additionally, preliminary experiments with open-source generators such as LLAMA2-13B led to weak candidate pools that worsened instruction-tuned models’ MD performance.
Instruction Format. We format the MD instruction inputs to place the generated instruction after the source documents, but alternate formats may be equally effective. Since the expected output length of downstream MD tasks has a high variance (e.g., keyword based QA vs. long-form summarization) we include a brief direction indicating the expected length of the answer (i.e., expected number of words or sentences).222 Our preliminary experiments demonstrated non-negligible performance gains on downstream MD tasks with the attachment of such length direction. This direction is appended to the end of the finalized instruction input, separated by a space (Fig. 3 in App. A.2).
Figure 2: Fine-grained scoring criteria utilized by MDCureRM for RM-as-a-Judge evaluation of candidate instruction quality in the Filtering phase of MDCure.
3.2 Filtering Phase
The generation prompts, while carefully designed, can still yield noisy or unsuitable instructions for MD settings. To address this, we apply a model-based filtering phase to remove low-quality, non-factual, or single-document instructions. Inspired by recent work on fine-grained RMs (Wu et al., 2023; Wang et al., 2024b, c), we train a fine-grained, MD-specific reward model, MDCureRM, to evaluate each instruction-answer pair based on six different criteria. These criteria capture both the overall quality of the instruction-response pairs and their effectiveness in handling multi-document content and are depicted in Fig. 2. This fine-grained RM-as-a-Judge mechanism refines the final curated instruction set to yield high-quality training data.
Training Setup. We adopt a multi-objective reward modeling framework (Wu et al., 2023; Wang et al., 2024b) to train MDCureRM. Traditionally, reward models for LLM alignment are trained with Bradley-Terry loss (Bradley and Terry, 1952) on pairwise data with binary preference annotations. However, we include the fine-grained information present in multi-objective ratings rather than binarizing the data, found to be effective in Wang et al. (2024c). Each training example for MDCureRM consists of a prompt, response, and a 6-dimensional rating vector, with each dimension corresponding to a specific reward objective. We apply multi-objective regression to learn from these ratings, utilizing a regression loss function.
Data. Training data for MDCureRM is obtained using an LLM-based pipeline without requiring human annotation. We first use GPT-4o-mini (OpenAI et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) to generate MD instruction-answer pairs of varying quality. [^3] We select these models due to their strong instruction-following capabilities, while considering the generation costs. This results in roughly 20,000 data points. We then prompt GPT-4o to assign scores for each sample according to the six criteria, to be utilized as annotations to train the reward model. The specific prompt used to do so can be found in App. A.5. The target reward scores are normalized to the range $[0,1]$. Although using GPT-4o directly as the RM is also possible, it is prohibitively costly to scale. Training MDCureRM ensures both cost-effectiveness and scalability for filtering large quantities of long-context data.
Implementation. MDCureRM utilizes the Llama3-8B architecture, with weights initialized from a Bradley-Terry reward model (Xiong et al., 2024) trained from Llama3-8B-Instruct (Dubey et al., 2024). The original output layer is replaced with a linear regression layer, outputting a 6-dimensional rating. The regression head processes the final-layer hidden states. We use Mean Squared Error (MSE) between target and predicted rewards as the loss function, freezing the base model during training, following Wang et al. (2024c).
Data Filtering. During inference, MDCureRM generates a 6-element rating for each candidate instruction. These scores are re-scaled to a 1-5 range and combined in a weighted average (details in App. A.6). The top $N$ highest-scoring samples are selected as the final instructions.
4 Experimental Setup
We conduct extensive experiments to evaluate the efficacy of MDCure for improving multi-document modeling capabilities of LLMs. In particular, we provide evidence for the following:
  • The scalability of MDCure as we vary the size of the generated instruction data and the size of the base LLM being instruction-tuned.
  • The importance of the Filtering stage for curating high-quality MD instruction data.
  • The superiority of MDCure’d models versus pre-trained MD and long-context baselines in 0-shot prompting settings.
  • The benefits of MDCure over existing synthetic MD-focused pre-training datasets.
  • The generalizability of MDCure to few-shot and long-context MD instruction data generation.
  • The value of different prompt approaches toward generating effective MD instructions.
In addition, we conduct ablations to demonstrate the criticality of each of the MDCureRM scoring criteria and the effect of their relative weightings. We discuss alternative architectures and the design of MDCureRM in App. B.2. We further provide a meta-evaluation to analyze the alignment of MDCureRM with human annotations versus GPT-3.5-Turbo in App. D, where we observe moderate agreements in line with LLM evaluation literature (Liu et al., 2024b). As we discuss later, we observe strong indication that MDCureRM provides consistent empirical improvements.
Models and Training Details. We apply MDCure on several LLMs across different widely used model families and sizes: FlanT5 (Chung et al., 2022), Qwen2 (Yang et al., 2024), and LLAMA3.1 (Dubey et al., 2024). We chose these models as they are capable open-source instruction-following models with varying architectural designs and parameter scales. We utilize the instruction-tuned variants as our base models to showcase the complementary effect of MDCure instructions versus typical IT data. To assess the performance of MDCure at scale, we create MDCure instruction datasets of size 12000, 36000, and 72000444Dataset sizes were determined as a function of API costs and based on the sizes of representative IT datasets used in related works (Bai et al., 2024; Köksal et al., 2023). and use them to fine-tune FlanT5-Base and -Large (250M & 750M), Qwen2-Instruct (1.5B & 7B), and LLAMA3.1-Instruct (8B & 70B). Hereafter, we refer to resulting models that have undergone MDCure IT as “MDCure’d” models. To assess the efficacy of MDCureRM, we additionally repeat these experiments using unfiltered and GPT-3.5-filtered instruction sets. [^5] We utilize GPT-3.5-Turbo for comparison as it is a more cost-effective alternative to commonly used scorer GPT4. Cheaper models such as GPT-4o-mini had not yet been released at the time of our starting experimentation. Further details regarding data and training setup are in App.s A and B.1.
Baselines. Beyond their base model counterparts, our MDCure’d models are compared against state-of-the-art pre-trained models for multi-document processing, namely PRIMERA and QAMDen, which are comparable in scale to FlanT5-Base and -Large, as well as several strong open-source long-context LLMs, namely LongAlign-7B (Bai et al., 2024) and ProLong-8B-64k-Instruct (Gao et al., 2024b), which are comparable to Qwen2-7B-Instruct and LLAMA3.1-8B-Instruct, and Jamba-1.5B-Mini (Labs, 2024) which has 54B parameters. We also compare MDCure’d models to SOTA general LLMs GPT-4o and Gemini 1.5 Pro (Team et al., 2024). Additionally, we inspect the effect of our instruction data versus baseline multi-document and long-context pre-training data at the 7B- and 8B-scales by fine-tuning Qwen2-7B-Instruct with PRIMERA and QAMDen data and ProLong-8B-64k-Instruct with MDCure data (72k). Further details regarding baselines are in App. C.2.
Benchmarks. We evaluate all baselines and MDCure’d models over a range of long- and short-form tasks through use of two challenging MD and long-context benchmarks: SEAM (Lior et al., 2024) and ZeroScrolls (Shaham et al., 2023). Within SEAM, we specifically consider the MultiNews, OpenAsp, MuSiQue, ECB+, and SciCo datasets to ensure coverage of a spread of domains and MD tasks. We additionally profile 0-shot prompting performance over a suite of four widely used multi-document benchmarks: WikiHop (Welbl et al., 2018) and HotpotQA-distractor (Yang et al., 2018) for MHQA, Multi-XScience (Lu et al., 2020) for MDS, and QMDSCNN (Pasunuru et al., 2021) for QMDS. While these datasets were built for use in fine-tuning evaluation, we draw upon the prompting approach proposed by Shaham et al. (2023) to formulate effective 0-shot prompts for each model. App.s C.1 and C.3 provide additional details.
Metrics. For SEAM and ZeroScrolls we employ the metrics and corresponding implementations defined by each for their component datasets. For WikiHop and HotpotQA-distractor, we use the standard F1 score. For Multi-XScience and QMDSCNN, we utilize LLM scores issued by GPT-3.5-Turbo and Gemini 1.5 Flash, respectively, to assess summarization coherence, relevance, fluency, and consistency. This follows recent adoption of LLM-as-a-Judge as a proxy for human-level rating (Liu et al., 2024c; Ye et al., 2024), given the limitations of traditional metrics like ROUGE for long-form generation quality assessment. (Sai et al., 2019; Gao et al., 2024a). Further details regarding the evaluator LLMs and the prompts are in App. C.4.
5 Results
IT Data Setting HQA WH MXS QC SEAM ZS Score Avg
ECB+ MN MSQ OA SC Avg
FlanT5 Base None 4.4 45.1 1.74 2.16 0.0 5.8 0.2 2.6 0.1 1.7 13.1 10.2
Unfiltered 31.7 47.5 4.04 2.35 0.0 6.7 0.1 2.7 0.2 1.9 21.0 16.1
GPT-Filtered 44.1 46.0 4.16 2.44 0.0 6.0 0.2 2.7 0.0 1.8 21.3 16.8
MDCureRM 47.3 48.3 4.22 2.58 0.0 7.5 0.3 2.7 0.0 2.1 22.6 17.9
Large None 24.4 54.6 3.19 2.78 1.1 7.9 0.1 3.1 0.3 2.5 23.2 17.3
Unfiltered 46.9 62.9 4.10 2.91 0.7 8.0 0.0 3.8 0.1 2.5 24.2 19.5
GPT-Filtered 46.3 65.1 4.13 2.88 0.5 8.3 0.0 3.9 0.1 2.6 24.2 19.7
MDCureRM 49.6 66.1 4.19 2.97 1.2 9.1 0.1 4.2 0.0 2.9 25.3 20.5
Qwen2-Ins 1.5B None 21.5 17.8 4.20 3.30 9.4 10.6 0.5 5.0 13.0 7.7 24.0 17.1
Unfiltered 32.9 30.5 4.24 3.57 15.7 12.0 0.4 5.0 16.8 10.0 23.9 18.9
GPT-Filtered 33.3 32.9 4.24 3.66 16.1 12.0 0.5 5.1 15.9 9.9 24.2 19.3
MDCureRM 37.7 34.8 4.25 3.73 16.4 12.0 0.6 5.9 18.1 10.6 25.6 20.5
7B None 30.5 39.6 4.30 3.57 5.0 11.9 0.5 6.4 13.1 7.4 23.9 18.6
Unfiltered 40.5 43.3 4.26 3.79 8.0 15.3 0.5 6.6 11.9 8.5 26.3 20.9
GPT-Filtered 42.0 44.0 4.26 3.84 8.7 15.3 0.9 6.6 12.1 8.7 28.7 22.3
MDCureRM 44.7 46.0 4.28 3.93 13.8 15.4 0.6 6.7 14.9 10.3 29.8 23.6
+PRIMERA 44.6 45.4 3.96 2.17 5.9 15.3 0.5 6.6 12.4 8.1 27.7 21.8
+QAMDen 42.6 45.8 3.92 2.37 9.5 15.3 0.5 6.4 12.4 8.8 28.5 22.3
Llama3.1-Ins 8B None 35.5 27.1 4.28 2.94 10.5 15.0 0.6 7.6 17.4 10.2 18.7 16.2
Unfiltered 37.6 34.4 3.81 4.07 15.3 16.3 0.5 7.8 18.9 11.8 28.6 22.3
GPT-Filtered 38.0 42.3 4.29 3.95 8.7 16.0 0.6 7.8 18.3 10.3 29.6 22.9
MDCureRM 44.7 43.7 4.29 4.22 16.3 16.4 0.6 7.9 18.5 11.9 30.9 24.5
70B None 53.9 38.1 4.28 3.97 25.1 21.9 0.6 6.1 11.5 13.0 36.4 27.9
Unfiltered 55.9 40.6 3.99 3.17 3.6 21.9 0.6 6.3 11.2 8.7 37.2 27.3
GPT-Filtered 57.4 41.2 3.97 3.37 4.9 22.0 0.7 6.4 10.7 8.9 37.4 27.6
MDCureRM 58.4 45.5 4.28 3.99 25.2 22.1 0.7 6.7 12.0 13.3 37.7 29.2
    Baselines PRIMERA 0.4 0.5 3.18 1.09 7.1 7.7 0.5 3.9 10.4 5.9 4.2 4.1
QAMDen 1.8 1.9 2.86 1.22 0.0 0.4 0.0 1.0 3.1 0.9 3.8 2.6
LongAlign-7B 10.4 14.3 4.15 3.75 11.5 16.5 0.0 4.1 16.8 9.8 23.2 16.5
ProLong-8B 43.6 34.3 3.86 2.46 9.5 17.4 0.4 10.7 18.0 11.2 32.2 24.3
Jamba 1.5 Mini 47.5 41.8 4.24 3.92 20.1 14.2 0.0 5.3 20.4 12.0 35.1 26.8
GPT-4o 57.5 50.0 4.50 4.25 8.9 18.6 0.3 14.2 28.9 14.1 39.8 30.8
Gemini 1.5 Pro 66.6 55.6 4.22 3.59 21.4 17.8 0.6 15.1 30.1 17.0 32.0 28.2
Table 1: Evaluation of MDCure’d models versus baselines on 6 benchmarks in the zero-shot prompting setting. Dataset abbreviations are described in App. C.1.1. Full results on ZeroScrolls are provided in App. E. Rows specified by MDCureRM refer to our full MDCure pipeline applied to the corresponding base model and size. Bold numbers show best performance in each group.
5.1 Main Results
We report experimental results in Table 1, with full results on ZeroScrolls included in App. E. For fair comparison, we report all MDCure results based on training with 72K samples, which led to best performance overall across all models. Our key findings are summarized as follows.
MDCure instructions are effective across model families and sizes. MDCure exhibits the best performance across all base models, demonstrating a clear and consistent advantage relative to the non-MDCure’d base models over all benchmarks. In particular, we see improvements of up to 975%, 96%, 143%, 434%, 39%, and 73% for HotpotQA, WikiHop, Multi-XScience, QMDSCNN, SEAM, and ZeroScrolls, respectively. MDCure’d LLAMA3.1-70B-Instruct achieves the best results across the 6 benchmarks, followed by MDCure’d Qwen2-7B-Instruct. As our models are trained from instruct-versions of base LLMs, it is clear that MDCure instruction data imparts complementary abilities versus typical IT data.
The relative improvement of MDCure’d models versus corresponding base models varies across model families, and improvements tend to shrink with model size. For FlanT5, MDCure provides 75.5% average improvement at the 250M-parameter scale (FlanT5-Base) versus 18.5% at the 750M-parameter scale (FlanT5-Large). For LLAMA3.1, MDCure provides 51.2% average improvement at the 8B scale versus 4.7% at the 70B scale. For Qwen2, we observe alternatively that MDCure gains shift from 19.9% average improvement at the 1.5B scale to 26.9% at the 7B scale. More generally, these results suggest the importance of instruction quality may weaken with base model scale and capability, similar to findings by Ivison et al. (2023).
Filtering with MDCureRM is essential to improve MD capabilities of LLMs. Comparing to models trained with unfiltered MDCure instruction samples, MDCure’d models which have been trained with the highest-scoring MDCureRM-filtered instructions surpass them across all benchmarks and base models. Furthermore, MDCureRM outperforms GPT-3.5-Turbo as a judge of MD instruction quality and efficacy, evidenced by the superior performance of MDCure’d models versus those trained with GPT-filtered instructions.
MDCure improves LLM performance on both multi-document tasks and single-document long-context tasks. Across MD benchmarks, MDCure enables models to achieve improved performance in all settings. In addition to MD tasks, MDCure generally boosts single-document long-context performance, particularly in tasks from ZeroScrolls, which emphasize general long-context capabilities. These trends are consistent across downstream task domains (e.g., science, literature), suggesting the MDCure procedure can boost MD performance in a domain-general way.
Model HQA WH MXS QC SEAM ZSS Avg
ProLong (PL) 43.6 34.3 3.86 2.46 11.2 32.2 24.3
PL+MDCure 45.8 45.1 3.93 3.40 12.5 32.2 25.4
Table 2: Our MDCure pipeline’s performance when applied to an already strong long-context LLM, ProLong-8B (Gao et al., 2024a).
Our best MDCure’d models consistently outperform pre-trained long-context baselines. For all benchmarks, our MDCure’d models achieve results that exceed the performance of size-comparable baselines [^6] We consider Jamba 1.5 Mini (54B) comparable with LLAMA3.1-70B-Instruct, as their parameter scales are within the same ballpark, falling within the same order of magnitude. by a substantial margin, confirming the efficacy of our design. For a more controlled setup, we take a strong open-source long-context LLM Prolong-8B (Gao et al., 2024a)) and fine-tune it using MDCure. The goal of this experiment is to study if our data provides additional advantages to long-context continual pre-training and fine-tuning. Table 2 demonstrates the results. Note that Prolong has been trained on large datasets totaling $\sim$81B tokens to improve long-context performance. As we observe, fine-tuning ProLong-8B with MDCure data which is much smaller ($\sim$162M tokens) yields further improved performance, indicating the complementary value of our MD training data to long-context training.
Our MDCure pipeline improves over existing multi-document synthetic data pipelines. We also compare our MDCure pipeline with existing MD-focused synthetic pre-training datasets inluding PRIMERA (Xiao et al., 2022) and QAMDen (Caciularu et al., 2023). As indicated by the “+PRIMERA” and “+QAMDen” rows in Table 1, MDCure data offers performance improvements across benchmarks at the 7B-scale. The weak performance of the original PRIMERA and QAMDen models (in the Baselines rows) can be explained as they are intended to serve as strong base pre-trained models for finetuning as opposed to zero-shot prompting settings. Overall, our results emphasize the utility of synthetic MD instructions as a promising approach for improving MD capabilities of LLMs.
5.2 Influence of Data Scale
# IT Samples HQA WH MXS QC SEAM ZSS
FlanT5-Base
None 4.4 45.1 1.74 2.16 1.7 13.1
12K RM-Filtered 43.8 48.4 4.16 2.43 1.9 22.0
36K RM-Filtered 45.7 46.1 4.17 2.55 1.9 22.2
72K RM-Filtered 47.3 48.3 4.22 2.58 2.1 22.6
72K Unfiltered 31.7 47.5 4.04 2.35 1.9 21.0
Qwen2-7B-Instruct
None 30.5 39.6 4.30 3.57 7.4 23.9
12K RM-Filtered 44.1 45.4 4.26 3.72 8.9 25.9
36K RM-Filtered 44.4 45.8 4.26 3.81 8.6 28.1
72K RM-Filtered 44.7 46.0 4.28 3.93 10.3 29.8
72K Unfiltered 40.5 43.3 4.26 3.70 8.5 26.3
Table 3: Performance at different training data scales.
In order to understand the importance of data quantity, we compare the impact of fine-tuning with MDCure instruction datasets of various sizes. We experiment with FlanT5-Base and Qwen2-7B-Instruct as the base models to demonstrate the effect of data size at different model scales. Results are compared versus the non-MDCure’d baselines and shown in Table 3. We observe consistent, albeit modest, improvements to downstream MD performance as we scale up the instruction data size. Notably, use of 12K MDCure-filtered samples consistently surpasses performance with use of 72K unfiltered samples for both base models, consistent with other work finding that use of only a few thousand high-quality instruction samples is sufficient for alignment (Zhou et al., 2023). This evidences MDCureRM’s ability to promote data efficiency and reduce costs involved in generating and using an effective MD instruction-tuning dataset.
5.3 Generalization to Few-Shot & Long-Context Training
We investigate the scalability of MDCure-generated instructions to the few-shot prompting and long-context settings. The few-shot setting is of interest as it provides an inexpensive alternative for adapting models to new tasks with limited compute (Dong et al., 2023). Likewise, long-context processing is important for real-world tasks over “extreme” numbers of documents such as extreme MDS (Lu et al., 2020). We extend MDCure instruction data to these settings either by appending distractor articles to existing MDCure instruction inputs or by packing few-shot exemplars into the input context up to a context length of 32k tokens. [^7] Context length selected based on GPU memory limits. Distractor articles are identified by embedding all NewSHead documents using the Sentence Transformers model all-distilroberta-v1, [^8] We choose this embedding model from https://www.sbert.net/docs/sentence_transformer/pretrained_models.html due to its strong performance, long context length, and compatibility with cosine similarity. and taking the most cosine-similar documents to the documents in the source cluster. We choose similar documents instead of performing random selection to increase the difficulty of the instruction task and encourage identification of relevant documents in a long context, and since recent work finds using sequences of related documents during training can assist models to read and reason across document boundaries (Shi et al., 2023). For few-shot training, exemplars are chosen randomly, with each exemplar appearing once across the duration of training.
HQA WH MXS QC SEAM ZSS
Zero-Shot Prompting Evaluation
LLAMA3.1-8B-Instruct 35.5 27.1 4.28 2.94 10.2 18.7
+MDCure 41.0 40.2 4.30 3.94 12.0 31.7
+MDCure +LongContext 38.9 40.5 4.31 3.97 12.1 32.5
5-Shot Prompting Evaluation
LLAMA3.1-8B-Instruct 57.5 44.2 3.97 2.93 15.1
+MDCure 63.9 47.0 4.04 3.32 17.3
+MDCure +Few-Shot 65.0 49.2 4.08 3.32 17.9
Table 4: Efficacy of MDCure for few-shot and long-context MD instruction data generation.
We conduct experiments with LLAMA3.1-8B-Instruct as the base model, with details of the training setup described in App. B.1. In the long-context setting, we use the same zero-shot prompting evaluations as before, comparing model performance when trained using distractor-extended instruction data versus when trained with non-extended instruction data and versus the non-MDCure’d baseline. In the few-shot setting, we perform similar comparisons but utilize 5-shot prompting during evaluation, leaving out ZeroScrolls since it is built for the zero-shot setting. Results are displayed in Table 4. We find that combining MDCure instruction data with longer contexts improves performance for most benchmarks. Similarly, incorporating multiple MDCure datapoints together to construct few-shot training inputs yields consistent improvements in few-shot evaluation. This confirms that our proposed instruction generation pipeline is extensible and beneficial for long-context tasks.
5.4 Effect of Generation Approach
The efficacy of MDCure is dependent on the quality of initial candidate instruction generations. To this end, we investigate the impact of several different prompting approaches on downstream instruction-tuned model performance. We consider only the zero-shot prompting setting given the high cost of prompting GPT models with few-shot examples that each contain multiple long documents.
In our preliminary investigation, we find that prompt templates emphasizing the MD nature of generated instructions tend to yield higher-scoring results (e.g., suggesting penalties if generated instructions concern only a single document). Thereafter, we explore the utility of providing general vs. highly detailed specifications for the style of the generated instruction-answer pairs. We ultimately converge on two classes of prompt templates which we refer to as either General or Style-Specific. Details are given in in App. A.2.
HQA WH MXS QC SEAM ZSS
FlanT5-Base
Baseline (No IT) 4.4 45.1 1.74 2.16 1.7 13.1
General 42.3 40.5 4.09 2.34 1.1 20.6
Style-Specific 37.9 48.1 4.13 2.20 1.3 20.0
Both 43.8 48.4 4.16 2.43 1.9 22.0
Qwen2-7B-Instruct
Baseline (No IT) 30.5 39.6 4.30 3.57 7.4 23.9
General 40.8 44.9 4.22 3.66 7.8 24.2
Style-Specific 40.2 45.8 4.24 3.63 7.8 23.2
Both 44.1 45.4 4.26 3.72 8.9 25.9
Table 5: Efficacy of different prompt approaches.
We examine the impact of General vs. Style-Specific prompt templates by fine-tuning FlanT5-Base and Qwen2-7B-Instruct using MDCureRM-filtered datasets created using templates from one or both categories. We limit the number of samples to 12K since §5.2 finds this quantity to be sufficient for strong improvements to MD task performance. When both templates are used, we employ a 1:3 ratio of General:Style-Specific instructions, found to be most effective in initial experiments.
Results are shown in Table 5. We observe that use of General or Style-Specific templates alone to produce instructions does not always lead to downstream MD performance improvements (e.g., FlanT5-Base IT’ed with General-based instructions does not surpass base model performance on WikiHop). In comparison, combining the two gives the best performance gains in a consistent fashion across benchmarks.
5.5 Ablations on Scoring Criteria
To justify our choice of scoring criteria in MDCureRM, we conduct an ablation study to show the advantage of combining all six criteria, rather than using a subset of them. We also evidence the benefit of applying greater weight to MD-specific criteria when computing the overall score for each candidate instruction, as opposed to computing an evenly weighted sum (details in App. A.6). Ablation experiments were run using FlanT5-Base and Qwen2-7B-Instruct. Table 6 shows the results. Across all benchmarks, filtering instruction data using a weighted average of six scoring criteria with emphasis on multi-document factors yields the best results. Further, removing any one criterion or utilizing even weighting across criteria leads to worsened IT’ed performance.
HQA WH MXS QC SEAM ZSS
FlanT5-Base
all criteria, MD emphasis 43.8 48.4 4.16 2.43 1.9 22.0
all criteria, evenly weighted 43.4 47.1 4.16 2.35 1.8 21.2
without Relevance 37.6 46.1 4.12 2.26 1.6 20.4
without Coh. & Fac. 40.1 46.0 4.14 2.35 1.5 20.4
without Creativity 38.4 44.2 4.15 2.33 1.3 20.8
without Context Int. 14.1 45.4 3.84 2.06 1.5 17.0
without Inter-Doc. Rel. 40.9 45.8 4.10 2.41 1.5 21.5
without Complexity 40.2 45.5 4.11 2.44 1.4 20.0
Qwen2-7B-Instruct
all criteria, MD emphasis 44.1 45.4 4.26 3.72 8.9 25.9
all criteria, evenly weighted 41.3 43.1 3.97 3.10 8.5 25.4
without Relevance 37.5 40.7 4.25 3.69 8.1 25.0
without Coh. & Fac. 41.9 43.8 4.27 3.63 8.4 23.2
without Creativity 43.6 44.3 4.23 3.72 8.9 25.4
without Context Int. 37.1 42.4 4.24 3.62 8.5 25.0
without Inter-Doc. Rel. 39.0 42.4 4.24 3.65 8.5 25.7
without Complexity 37.1 41.8 4.25 3.65 8.5 25.7
Table 6: Scoring criteria ablation study results.
6 Conclusion
In this work, we presented MDCure, a procedure for automatically generating high-quality multi-document instruction tuning data to improve any base LLM’s ability to leverage long range dependencies for MD tasks. MDCure addresses the challenge of data scarcity in MD task domains and is general and scalable across model families and sizes. We further introduced MDCureRM, a novel RM-as-a-Judge LLM that can be used to filter MD instruction data more cheaply and effectively than GPT-3.5-Turbo. Our extensive experiments on various LLMs in multiple task domains demonstrate that MDCure adds value beyond typical instruction tuning and can substantially boost MD performance while overcoming limitations of previous pre-training-based approaches. Future work could explore extending MDCure to other topic domains, or use of MDCureRM to train a custom MD instruction generator model.
7 Limitations
While we have obtained promising experimental results, we recognize several limitations of this work. In our experiments we limit generation of MD instruction data to the news domain to preserve comparability to prior state-of-the-art MD pre-training approaches. Applying MDCure to document sets in other domains and text genres (e.g., science, mathematics) may yield further boosts to instruction diversity and generalization performance of resulting IT’ed models. While MDCure is able to accommodate use of arbitrarily large document collections for instruction generation, our experiments are limited to generation over 3-5 articles due to limited cluster size in the NewSHead source dataset. Scaling up the number or length of context documents may require use of more powerful and costly generator LLMs such as GPT4 to ensure sufficient quality of the raw generations, which place an upper bound on the overall composition of any possible resulting instruction set. This may limit the scalability of the approach; however, combined with the costs saved through use of our custom RM-as-a-Judge model, we believe MDCure provides a promising avenue to more effectively balance costs associated with creating high-quality instruction data versus existing synthetic data pipelines.
Experimentally, our training is limited to 4096 context length due to compute restrictions. Given the results in §5.2 and §5.3, we expect that our MDCure’d models could achieve greater performance on MD benchmarks if more instruction samples and greater training context length were used. The design and application of our approach is limited to texts in English, suggesting limited utility of MDCure for non-English tasks. Further, despite our exploration of a wide range of prompts for raw instruction generation, it is plausible that slight changes in phrasing could still affect the quality of resulting generations, and different generator models may need different prompts to achieve similar effect. Potential improvements may arise from generating instructions and answers separately as opposed to together (Zhao et al., 2024a), and from incorporation of short-context data during instruction-tuning (Bai et al., 2024; Gao et al., 2024b). Nonetheless, instruction complexity and answer correctness are constrained by the ability for the generator model to perform such MD tasks.
8 Ethics Statement
As MDCure creates instruction data based on real-world text corpora, one potential consequence is amplification of biases from the source texts in resulting IT’ed LLMs. Source texts are generally not filtered prior to use in MDCure, so it is possible to enable creation of instruction tuning samples concerning unethical or sensitive content. As with any use of LLMs, there is a need to ensure safety of IT’ed models’ responses through incorporation of red-teaming or other safety measures. Use of our factuality criterion in assessing instruction data reduces the chances for MDCure to lead to unfactual examples due to hallucination during instruction generation. However, faithfulness to false anecdotal information in source texts could take precedence, leading to propagation of potentially biased content. More beneficially, we anticipate our work will improve LLMs’ efficacy in processing long documents or large document collections, which can accelerate value generation in domains such as business analytics and scientific literature mining.
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2139841, We are grateful for the compute support provided by Google. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or Google.
Appendix A Dataset Construction
A.1 Source Texts
As noted in §3.1, our experiments base data generation on the NewSHead corpus (Gu et al., 2020). NewSHead consists of 369940 clusters of 3-5 related English news articles curated for the task of news story headline generation. We utilize the already-preprocessed version of the corpus supplied by Caciularu et al. (2023) at https://storage.googleapis.com/primer_summ/processed_pre-training_data.tar.gz. This follows the pre-processing procedure described in Xiao et al. (2022).
A.2 Data Generation Prompts
During the Generation phase of MDCure, we employ two types of prompt templates to encourage the generator model to produce a diverse set of instruction data. We distinguish these templates as either General or Style-Specific.
General templates are inspired by Köksal et al. (2023) and broadly ask the generator model to produce an instruction-answer pair adhering to a certain answer length (e.g., 1-2 words, 5 or more sentences). Emphasis is placed on creation of instructions that cannot be answered if any one context document is removed, and we employ various phrasings of this specification. We utilize a total of 14 general prompts. General templates A-D (Figure 4) are inspired by recent work that demonstrates the efficacy of grounding synthetic generations in pre-selected answer content from a given source. Modeling after (Köksal et al., 2023), for these 4 prompt templates, we consider all possible combinations of two related documents in NewSHead and split the contents of every possible document pair into texts 1-3 sentences in length. Each segment is embedded via HuggingFace Sentence Transformers (Reimers and Gurevych, 2019) and aligned with the others via cosine similarity. Distinct segments are then paired according to highest alignment for according use in the prompt templates. To ensure segments are not too similar, we set a maximum similarity threshold of 0.70 per pair. General templates E-N (Figures 5 and 6) consider 2 or more context documents per generation.
On the other hand, Style-Specific templates (Figure 7) are adapted from Zhu et al. (2024) and employ a single phrasing while incorporating varying combinations of constraints regarding instruction complexity (e.g., analysis, recall), type (e.g., paraphrasing, inference), style (e.g., imperative, interrogative), and output length (e.g., 3-4 words, 3 sentences), with pre-defined options for each specification category. The aim is to enforce constraints on diversity of the generated candidate data. As defined below, we consider 4 characteristics, 4 types, and 3 styles of instruction, in combination with 8 possible answer lengths (Figure 8). Each possible resulting prompt template considers 2 or more context documents per generation.
To ensure a balanced composition for the pool of candidate instructions, we sample uniformly from all possible prompt templates during generation for each cluster of related documents. As mentioned in §3.1, once all instruction-answer pairs are generated, we append a brief direction indicating the expected length of the answer to the end of each instruction input (Figure 3). Options for this direction differ depending on whether an instruction was created using a General or Style-Specific template (Figure 9).
Figure 3: Illustrative example of input format.
General Template A Snippets: ‘{segment1}’, ‘{segment2}’ Context Paragraphs: ‘{doc1}’, ‘{doc2}’ Based on the given snippets and context paragraphs, construct an instruction-answer pair such that (1) the answer is based on the two snippets and (2) the instruction is a plausible prompt or question to which the answer would be the expected response. Make sure both snippets are required to answer the instruction. You will be penalized if the instruction concerns only one snippet. Format your response as: Instruction: <prompt or question> Answer: <answer>
General Template B Snippets: ‘{segment1}’, ‘{segment2}’ Based on the given snippets, construct an instruction-answer pair such that (1) the answer is yes and (2) the instruction is a plausible prompt or question to which yes would be the expected response. Make sure the answer does not conflict with the information in the snippets. You will be penalized if the instruction-answer pair is unfactual. Do NOT mention the snippets in the instruction. Format your response as: Instruction: <prompt or question> Answer: <yes>
General Template C Snippets: ‘{segment1}’, ‘{segment2}’ Based on the given snippets, construct an instruction-answer pair such that (1) the answer is no and (2) the instruction is a plausible prompt or question to which no would be the expected response. Make sure the answer does not conflict with the information in the snippets. You will be penalized if the instruction-answer pair is unfactual. Do NOT mention the snippets in the instruction. Format your response as: Instruction: <prompt or question> Answer: <no>
General Template D Snippets: ‘{segment1}’, ‘{segment2}’ Context Paragraphs: ‘{doc1}’, ‘{doc2}’ Based on the given snippets and context paragraphs, construct an instruction-answer pair such that (1) the answer is a brief phrase and NOT a sentence and (2) the instruction is a plausible prompt or question to which the answer is the expected response. Make sure both snippets are required to answer the instruction. You will be penalized if the instruction concerns only one snippet. Make sure the answer is a brief phrase less than 7 words in length, with NO periods. You will be penalized if the answer is longer than 7 words or if the answer is a sentence. Format your response as: Instruction: <prompt or question> Answer: <answer>
Figure 4: General Prompt Templates A-D
General Template E Context Paragraphs: ‘{doc1}’, ‘{doc2}’ Based on the two given context paragraphs, construct an instruction-answer pair such that (1) the answer is summary of the two paragraphs and (2) the instruction is a plausible prompt or question to which the answer is the expected response. Make sure both paragraphs are required to answer the instruction. You will be penalized if the instruction concerns only one paragraph. Make sure the answer does not conflict with the information in the paragraphs. You will be penalized if the instruction-answer pair is unfactual. Make sure the answer is at least 5 sentences in length. Do not mention the context paragraphs in the instruction. Format your response as: Instruction: <prompt or question> Answer: <answer>
General Template F Context Paragraphs: ‘{doc1}’, ‘{doc2}’ Based on the two given context paragraphs, construct an instruction-answer pair such that (1) the answer is summary of the two paragraphs and (2) the instruction is a plausible prompt or question to which the answer is the expected response. Make sure both paragraphs are required to answer the instruction. You will be penalized if the instruction concerns only one paragraph. Make sure the answer does not conflict with the information in the paragraphs. You will be penalized if the instruction-answer pair is unfactual. Make sure the answer is less than 5 sentences in length. Do not mention the context paragraphs in the instruction. Format your response as: Instruction: <prompt or question> Answer: <answer>
General Template G {context_docs} A question or command that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed is: Question: <respond here> Answer: <respond here briefly>
General Template H {context_docs} What is a question or command that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed? Question: <respond here> Answer: <respond here briefly>
General Template I Articles: {context_docs} What is an exam question that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed? Exam Question: <respond here> Answer: <respond here briefly>
Figure 5: General Prompt Templates E-I
General Template J {context_docs} What is a question or command that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed? Question: <respond here> Answer: <respond here, feel free to use a single word or phrase>
General Template K {context_docs} A question or command that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed is: Question: <respond here> Answer: <respond here>
General Template L {context_docs} What is a question or command that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed? Question: <respond here> Answer: <respond here, using ONLY a single word or phrase>
General Template M Articles: {context_docs} Contrasting Question: <respond here> Answer: <respond here briefly>
General Template N Articles: {context_docs} What is an exam question that can ONLY be answered by utilizing ALL of the above documents and that CANNOT be answered if any one document is removed? Exam Question: <respond here> Answer Choices: <respond here> Answer: <answer letter only>
Figure 6: General Prompt Templates J-N
Style-Specific Template You’re proficient in crafting complex questions. Generate only one question and one answer that adheres to the provided #Articles#. Make sure the question and answer are factually consistent with the #Articles#. The question should meet the following criteria: 0. The person answering the question cannot see the #Articles#, so the question must not contain phrases like ‘Given the information provided’, ‘Based on the provided information’, or similar expressions that imply direct citations or references from #Articles#. 1. The question must REQUIRE synthesis of information in at least 2 of the provided documents in order to answer correctly. The more documents are involved the better. Ideally all documents are required to answer the question, such that losing any one of them will lead person answering the question to provide an incorrect response. You will lose your job if this criterion is not satisfied. 2. {complexity} 3. {type} 4. {style}. 5. It requires {answer_length} to answer correctly. The answer must be {answer_length} in length. ### Articles: {context_docs} Question: <respond here> Answer: <respond here>
Figure 7: Style-Specific Prompt Template.
Options for {complexity} Specification "It should be complex and require multiple-step reasoning across the documents to solve." "It demands critical thinking skills to analyze, evaluate, and synthesize multiple pieces of information from the different documents." "It demands integrating knowledge from multiple documents to address its multifaceted nature." "It should be simple and require only a few words to answer, yet utilize supporting evidence from at least 2 documents."
Options for {type} Specification "It is a Natural language inference question: Assessing if evidence supports a conclusion." "It is a Paraphrasing question: Rewording a statement while retaining its meaning." "It is a Summarization question: Condensing key information from a larger text." "It is an Informational question: Locating a specific piece of information in the given evidence."
Options for {style} Specification "It should be in the style of a command or imperative. For example, ‘Write a paragraph about…’ or ‘Describe the…’" "It should be in the style of a question or interrogative. For example, ‘What is the..?’ or ‘How do you…?’" "It should be in the style of a short phrase that serves as a query. For example, ‘today’s forecast.’ or ‘Donna’s car accident.’"
Options for {answer_length} Specification "1-2 words" "3-4 words" "a phrase of at least 5-6 words" "1-2 sentences" "3-4 sentences" "6 sentences" "8 sentences" "10 sentences"
Figure 8: Detailed options for Style-Specific prompt template.
Length Direction Options (General) "Answer briefly in 1-2 sentences." "Answer ‘yes’ or ‘no’" "Answer ‘yes’ or ‘no’" "Answer with a single word or brief phrase." "Answer with at least 5 sentences." "Answer with at most 5 sentences."
Length Direction Options (Style-Specific) "Answer with {answer_length}." "Answer using {answer_length}." "Respond with {answer_length}." "Respond using {answer_length}." "Formulate your answer in {answer_length}." "Reply with a {answer_length} answer." "Craft your response in {answer_length}." "Give a response that is {answer_length}." "Answer in around {answer_length}."
Figure 9: Options for direction regarding expected output length for instructions generated via General and Style-Specific templates.
A.3 Data Generation Hyperparameters
To elicit high-quality candidate instructions in a cost-effective manner for our experiments, we utilize GPT-3.5-Turbo as the generator model in MDCure with all inference hyperparameters set to their default values. The cost is approximately \$25.00 per 35,000 candidate instruction generations.
A.4 GPT Scoring Prompt
We use the prompt shown in Figure 10 to elicit effective scores from GPT-3.5-Turbo for the candidate generations in a zero-shot fashion. This prompt is adapted from that used in the self-curation step of Li et al. (2024); other prompt variations (e.g., additive scoring as in (Yuan et al., 2024), rubric-free scoring as in (Li et al., 2023b)) yielded relatively worse MD performance in models instruction-tuned with the resulting filtered samples and were therefore discarded from consideration.
Instruction: {generation_prompt} Response: {generated_instruction_answer_pair} Above are an Instruction from a user and a candidate Response from an AI Assistant. The goal of this AI Assistant is to generate a Response that effectively addresses the user’s Instruction and that, in order to answer, requires the ability to reason over multiple documents. The Response should be a targeted question, instruction, prompt, or task that requires the use of information from different positions in the provided texts. Evaluate whether the Response is a good example of how an AI Assistant should respond to the user’s Instruction. Assign a score to the Response using the following 5-point scale: 1: It means the Response is incomplete, vague, off-topic, or not exactly what the user asked for in the Instruction. Perhaps it provides an incomplete prompt. Or it can be answered without looking at source documents provided in the Instruction. Or some content seems missing, the opening sentence repeats user’s question, or it contains is irrelevant to the source documents or snippets provided in the Instruction. 2: It means the Response addresses some of the asks from the user but does not directly address the user’s Instruction. For example, the Response only leverages one of several source documents or snippets provided in the Instruction. Or the Response can be answered using only ONE source document or snippet, and thus does not effectively assess and require use of multi-document reasoning capabilities. 3: It means the Response is fair and addresses all the basic asks from the user. It is complete and self contained and is relevant to most of the documents or snippets provided, but not all. It may be somewhat helpful toward assessing an agent’s multi-document reasoning capability but still has room for improvement. 4: It means the Response is good quality. Specifically, the Response can only be answered by performing reasoning across most of the documents or snippets provided in the Instruction. The provided documents or snippets include all the information required to answer the Response, i.e. no information beyond that provided in the Instruction is needed. The Response has minor room for improvement, e.g. more concise and focused. 5: It means the Response is perfect, i.e. it can only be answered with strong ability to extract and synthesize information across the documents or snippets provided in the Instruction. The Response utilizes ALL documents or snippets provided in the instruction. The provided documents or snippets include all the information required to answer the Response. It is well-written and effective toward the AI Assistant’s goal and has no irrelevant content. Assess the Response and assign a rating score using this scale. Respond with "Score: <rating>".
Figure 10: Prompt used for GPT-3.5-Turbo to evaluate the quality of candidate MD instruction-answer pairs.
A.5 MDCureRM Training Data Prompt
As described in §3.2, we use the prompt shown in Figure 11 to generate fine-grained reward training data for MDCureRM.
Instruction Quality Rating Task Rate the quality of the generated instruction based on the provided documents, using a scale of 1-5. Only provide numerical scores, without any rationale or explanation. Relevance: Does the instruction align well with the content of the documents? Does it make sense given the provided information? Coherence & Factuality: Is the instruction-answer pair coherent, logical, and factually accurate? Does the answer appropriately address the instruction and is it well-supported by the documents? Creativity: How diverse or creative is the instruction in terms of question type (e.g., factual, inferential) and format (e.g., multiple-choice, open-ended)? Context Integration: How well does the instruction leverage and synthesize information from multiple documents to form a comprehensive response? Inter-Document Relationships: Does the instruction encourage understanding relationships (e.g., comparisons, contrasts, discrepancies) between different documents? Complexity: Does the instruction appropriately challenge the answerer to think critically and synthesize information from multiple sources? Input: Context: {context_docs} {instruction_sample} Output (provide only numerical scores, no rationale): Relevance: [score] Coherence & Factuality: [score] Creativity: [score] Context Integration: [score] Inter-Document Relationships: [score] Complexity: [score]
Figure 11: Prompt to elicit fine-grained score training data for MDCureRM.
A.6 MDCure Data Scoring
Given a candidate instruction-answer pair, MDCureRM generates a vector of six floating point values in the range $[0,1]$. To re-scale each value to the 1-5 range, we multiply each criterion score by 4 and add 1. To obtain the overall score for the instruction sample, we take a weighted sum of the resulting values, using a weight of $\frac{1}{9}$ for the general quality criteria (Relevance, Coherence & Factuality, Creativity) and a weight of $\frac{2}{9}$ for the multi-document specific criteria (Context Integration, Inter-Document Relationships, Complexity). In §5.5, we compare this weighting assignment with the evenly weighted alternative, which utilizes a weight of $\frac{1}{6}$ for each criterion in the case where all 6 criteria are used, and a weight of $\frac{1}{5}$ for each criterion in the case where one criterion is ablated at a time.
A.7 MDCure Dataset Statistics
We summarize statistics for the unfiltered, GPT-3.5-Turbo-filtered, and MDCureRM-Filtered instruction data as shown in Table 7. Quantitatively, we observe that the MDCureRM-Filtered data tends to have a slightly higher number of documents per sample compared to the GPT-filtered data. Meanwhile, the instruction lengths in the MDCureRM-Filtered and unfiltered pools are quite similar, both having longer and more detailed instructions on average, whereas the GPT-filtered pool features shorter instruction lengths. Complexity in terms of context length and instruction lengths is apparent across all data settings, but the MDCureRM-Filtered data handles the most complex tasks while GPT-filtered data leans towards handling smaller, more straightforward tasks.
Data Setting # Instr. Avg. # Context Docs Avg. Instr. Length
Unfiltered 221,835 3.4 1,754
GPT-Filtered 70,598 3.2 1,377
RM-Filtered 125,994 3.4 1,719
Table 7: Data statistics according to instruction composition.
Appendix B Training Details
B.1 Instruction Tuning
All FlanT5 models are trained using 8 NVIDIA A6000 GPUs or 8 A100 40GB GPUs. For LLAMA3.1 and Qwen2 fine-tuning experiments, we employ the widely-used Unsloth repository with HuggingFace to perform 4-bit quantization and Low-Rank Adaptation (LoRA). We choose to use LoRA over full fine-tuning due to compute limitations and since preliminary experiments yielded similar performance in both settings. LoRA adapters are used for every linear layer for all query-, key-, value-, output-, gate-, up-, and down-projection matrices with a rank of 16, alpha parameter of 32, and zero adapter dropout. As Unsloth only accommodates single-GPU training, all LLAMA3.1 and Qwen2 models are trained on a single A6000 or A100 GPU. All fine-tuning experiments utilized gradient checkpointing aside from those involving FlanT5-Base.
As the FlanT5 models could be trained with a maximum length of 4096 tokens without GPU memory overflow, we set the maximum input length of the training data to 4096 for all fine-tuning experiments beyond those in §5.3 to ensure comparability. Any samples exceeding this length were truncated from the right, by shortening component context documents to equal lengths and preserving the entirety of the subsequent instruction. For fine-tuning experiments with long-context and few-shot MDCure data (§5.3), we set the maximum input length to 32000 tokens.
All fine-tuning experiments utilized a global batch size of 64, AdamW optimization with $\beta_{1}=0.9$ and $\beta_{2}=0.98$, a linear learning rate schedule, warmup ratio of 0.15, dropout of 0.05, weight decay of 0.01, and the best of either 1e-4 or 5e-5 for the maximum learning rate. Fine-tuning is performed for 1100, 1650, and 2250 steps for the 12K-, 36K-, and 72K-training sample settings. For the long-context and few-shot training settings, we use a total of 36K samples, with the first 90% of training steps utilizing 4096 maximum input length and the remaining 10% utilizing 32000 maximum input length, in a similar fashion to Dubey et al. (2024). In terms of training data composition, unless otherwise stated, all models employed a 1:3 ratio of General:Style-Specific prompt template generations. Using the hardware specified above, all fine-tuning experiments require 24-120 GPU-hours.
In all settings, we determine the best checkpoint for each model by comparing checkpoint performances in the zero-shot prompting setting on 500 test samples from MultiNews, Multi-XScience, HotpotQA-distractor, WikiHop, and QMDSCNN; we use LLM evaluation to assess MDS performance and F1 and EM scores to assess MHQA performance. Evaluation details are provided in App. C.4. For language modeling settings, we apply the loss to the instruction input as well during training, per recommendation by Shi et al. (2024).
B.2 MDCureRM Training
To train MDCure RM, we utilize 8 A100 40GB GPUs. Training occurs for 4 epochs using a training set of size approximately 17K samples and validation and test sets of size approximately 1.5K. We set the global batch size to 64 and use AdamW 8-bit optimization, a linear learning rate schedule, warmup ratio of 0.1. We perform a small hyperparameter search over learning rates of 7e-6, 1e-5, 5e-5, 1e-4, 7e-4, and 1e-3.
Beyond use of a single regression head to construct MDCureRM as in Wang et al. (2024c), we investigate the impact of several variations in the architecture and design of MDCureRM. These variations include the addition of 2 or 3 linear layers to the backbone LLM as opposed to a single regression head, as well as use of an MLP head with attention pooling. Due to compute limitations we were unable to investigate the impact of unfreezing the backbone LLAMA3 model during MDCureRM training. Across these variations, we determined the best scoring model based on test loss, finding use of a single regression head and learning rate 1e-3 most effective. Small-scale experiments on FlanT5-Base demonstrated this improves over the MLP variant.
Appendix C Evaluation Details
C.1 Benchmark Descriptions
We provide details on the datasets used to benchmark our instruction-tuned models versus baselines. All benchmarks are in English.
HotpotQA (Yang et al., 2018) is a multi-hop QA dataset consisting of approximately 113,000 questions and paragraphs sourced from Wikipedia, where for each sample there are eight paragraphs that serve as distractors and two paragraphs containing the information necessary to answer the given question. The objective is to identify the correct answer and evidence spans in the text.
WikiHop (Welbl et al., 2018) is a multi-hop QA dataset consisting of approximately 43,000 samples constructed from entities and relations from WikiData and supporting documents from WikiReading. Each sample provides a question, 2-79 potential answers, and 3-63 supporting contexts. The aim is to arrive at the correct answer based on the context and question.
Multi-XScience (Lu et al., 2020) is a MDS dataset sourced from Arxiv and Microsoft scientific graphs. The aim is to produce the related works section for a query publication given the abstract of the query paper and several abstracts of other referenced works. Multi-XScience is considered relatively more difficult versus datasets such as Multi-News since it is less susceptible to positional and extractive biases.
QMDSCNN (Pasunuru et al., 2021) is a query-focused MDS dataset of approximately 287K samples created by extending the CNN/DailyMail single-document summarization dataset to the multi-document setting. Given paragraph-scale chunks of various news articles, the aim is to return relevant context portions given queried article titles.
SEAM (Lior et al., 2024) is a recent multi-document benchmark that addresses the sensitivity of LMs to prompt variations through repeated evaluations with controlled input-output formats on a diverse set of multi-document datasets. Tasks include multi-document summarization, multi-hop question-answering, and cross-document coreference resolution.
ZeroSCROLLS (Shaham et al., 2023) is a collection of datasets for long-context natural language tasks across domains of summarization, QA, sentiment classification, and information reordering. Successful performance on the benchmark requires synthesizing information across long texts.
C.1.1 Dataset Abbreviations
We provide a list of dataset name abbreviations in Table 8.
Dataset Name Abbreviation
HotpotQA-distractor HQA
WikiHop WH
Multi-XScience MXS
QMDSCNN QC
SEAM
MultiNews MN
OpenAsp OA
MuSiQue MSQ
SciCo SC
ECB+ ECB+
ZeroScrolls
GovReport GvRp
SummScreenFD SSFD
QMSum QMsm
NarrativeQA NrQA
Qasper Qspr
QuALITY QLTY
SQuALITY SQLT
MuSiQue MuSQ
SpaceDigest SpDg
BookSumSort BkSS
ZeroScrolls (ZS) Score ZSS
Table 8: Dataset name abbreviations used for results tables in the main text.
C.2 Baselines
PRIMERA (Xiao et al., 2022) is a 447M-parameter model initialized from LED-large (Beltagy et al., 2020) and trained using an entity salience-based pre-training objective over the NewSHead dataset. It previously stood as the state-of-the-art in multi-document summarization.
QAMDen (Caciularu et al., 2023) is a 447-parameter model initialized in PRIMERA sparse attention style over LED-large and trained using a QA-based pre-training objective over the NewSHead dataset. It surpasses PRIMERA in performance on MDS, QMDS, and MDQA tasks.
LongAlign (Bai et al., 2024) is a recipe for long-context alignment that enables improved long-context performance of fine-tuned models through a combination of data and packing strategies during training. It provides a series of instruction-tuned long-context models up to 13B parameters in size adapted from the ChatGLM3 (GLM et al., 2024) and LLAMA2 (Touvron et al., 2023) model families. We use LongAlign-7B for comparability to Qwen2-7B-Instruct and LLAMA3.1-8B-Instruct, which we use for MDCure experiments.
ProLong-8B-64k-Instruct (Gao et al., 2024b) is continued-trained and supervised-fine-tuned from the LLAMA-3-8B model, with a maximum context window of 64K tokens. It achieves state-of-the-art long-context performance versus similarly sized models and outperforms LLAMA3.1-8B-Instruct on many long-context tasks.
Jamba 1.5 Mini (Labs, 2024) is a speed-optimized model with 52B total parameters and 12B active parameters that is specifically trained for improved long-context handling and quality. It utilizes a hybrid transformer architecture.
C.3 Zero-Shot Evaluation Prompts
To perform zero-shot prompting evaluations on Multi-XScience, QMDSCNN, HotpotQA, and WikiHop, we devise effective zero-shot prompts for each dataset in a similar fashion to Shaham et al. (2023) and utilize the same input format as Shaham et al. (2023) as shown in Figure 12. For each model family (FlanT5, LLAMA3.1, Qwen2), we choose the best one of 5 tailored prompt variations based on base model performance to utilize for evaluations of all subsequent instruction-tuned models. We display the final prompts utilized for each model family in Figure 13. During inference, we use greedy decoding and limit the model generation length to at most the 90th percentile of the maximum target output length. The input context is truncated to length 8192 for every model, and for all results reported in the main text, we utilize 1500 test samples per dataset as opposed to using the entire test set for each dataset to accommodate cost limitations imposed by LLM evaluations for MDS.
Zero-Shot Input Format for HotpotQA {prompt} Question: {question} Supporting Documents: {supporting_documents} Answer:
Zero-Shot Input Format for WikiHop {prompt} Documents: {supporting_documents} Question: {question} Answer Candidates: {answer_choices} Answer:
Zero-Shot Input Format for Multi-XScience {prompt} Documents: {source_and_reference_abstracts} Related Work Section:
Zero-Shot Input Format for QMDSCNN {prompt} Query: {question} Articles: {articles} Summary:
Figure 12: Task input formats used for zero-shot prompting evaluation on HotpotQA, WikiHop, Multi-XScience, and QMDSCNN, adapted from Shaham et al. (2023).
Final Task Prompts for FlanT5 Models HotpotQA: "You are given a question and multiple supporting documents separated by ‘|||||’. Answer the question as concisely as you can, using a single phrase if possible." WikiHop: "You are given multiple supporting documents separated by ‘|||||’, a question, and list of answer candidates. Answer the question by selecting one of the provided answer candidates." Multi-XScience: "You are given several scientific documents, separated by ‘|||||’. Write the related-work section of a paper based on the given abstracts and articles. Answer in a single paragraph." QMDSCNN: "You are given a query and an article. Answer the query as concisely as you can by summarizing the relevant information from the article. Respond in a single paragraph."
Final Task Prompts for LLAMA3.1 Models HotpotQA: "You are given a question and several supporting documents. Answer the question as concisely as you can, using a single word or phrase." WikiHop: "You are given multiple supporting documents separated by ‘|||||’, a question, and list of answer candidates. Answer the question by selecting one of the provided answer candidates." Multi-XScience: "You are given several scientific abstracts. Write a related works section based on the abstracts. Answer in a short paragraph. Be very brief." QMDSCNN: "You are given a query and several articles. Answer the query as concisely as you can by supplying the relevant information from the articles. Respond briefly, in less than one paragraph."
Final Task Prompts for Qwen2 Models HotpotQA: "You are given a question and multiple supporting documents separated by ‘|||||’. Answer the question as concisely as you can, using a single word or phrase." WikiHop: "You are given multiple supporting documents separated by ‘|||||’, a question, and list of answer candidates. Answer the question by selecting one of the provided answer candidates." Multi-XScience: "You are given several scientific abstracts, separated by ‘|||||’. Write the related works section of a paper based on the given abstracts. Answer in a single paragraph." QMDSCNN: "You are given a query and several articles. Answer the query as concisely as you can by supplying the relevant information from the articles. Respond briefly, in less than one paragraph."
Figure 13: Task prompts yielding the best performance for base models in the FlanT5, LLAMA3.1, and Qwen2 model families, selected from among 5 tailored variations.
C.4 LLM Evaluation Details
As noted in §4, to evaluate MDS performance, we turn to LLM evaluation to overcome limitations of $n$-gram-based methods such as ROUGE in long-context settings (Shaham et al., 2023). We employ the open-source LLM evaluation framework G-Eval (Liu et al., 2023), which is one of the most highly human-correlated evaluation frameworks to date when paired with GPT-4. For a given document and summary, G-Eval determines the quality of the summary by issuing four fine-grained scores for Relevance, Coherence, Consistency, and Fluency. The overall LLM score for a given summarization response is obtained by taking the average of the four scores. We adapt G-Eval to the multi-document setting using the prompts shown in Figures 14, 15, 16, and 17.
G-Eval Prompt Template for LLM Evaluation of MDS Relevance You will be given one summary written for a set of related documents. Your task is to rate the summary on one metric. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Relevance (1-5) - selection of important content from the source. The summary should include only important information from the source documents. Annotators were instructed to penalize summaries which contained redundancies and excess information. Evaluation Steps: 1. Read the summary and the source documents carefully. 2. Compare the summary to the source documents and identify the main points of the articles. 3. Assess how well the summary covers the main points of the source documents, and how much irrelevant or redundant information it contains. 4. Assign a relevance score from 1 to 5. Example: Source Text: {documents} Summary: {summary} Evaluation Form (scores ONLY): - {Relevance}
Figure 14: G-Eval prompt for NLG evaluation of the relevance of a given summarization response.
G-Eval Prompt Template for LLM Evaluation of MDS Coherence You will be given one summary written for a set of related documents. Your task is to rate the summary on one metric. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Coherence (1-5) - the collective quality of all sentences. We align this dimension with the DUC quality question of structure and coherence whereby "the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to a coherent body of information about a topic. Evaluation Steps: 1. Read the source documents carefully and identify the main topic and key points. 2. Read the summary and compare it to the source documents. Check if the summary covers the main topic and key points of the source documents, and if it presents them in a clear and logical order. 3. Assign a score for coherence on a scale of 1 to 5, where 1 is the lowest and 5 is the highest based on the Evaluation Criteria. Example: Source Text: {documents} Summary: {summary} Evaluation Form (scores ONLY): - {Coherence}
Figure 15: G-Eval prompt for NLG evaluation of the coherence of a given summarization response.
G-Eval Prompt Template for LLM Evaluation of MDS Consistency You will be given one summary written for a set of related documents. Your task is to rate the summary on one metric. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Consistency (1-5) - the factual alignment between the summary and the summarized sources. A factually consistent summary contains only statements that are entailed by the source documents. Annotators were also asked to penalize summaries that contained hallucinated facts. Evaluation Steps: 1. Read the source documents carefully and identify the main facts and details they present. 2. Read the summary and compare it to the source documents. Check if the summary contains any factual errors that are not supported by the source documents. 3. Assign a score for consistency based on the Evaluation Criteria. Example: Source Text: {documents} Summary: {summary} Evaluation Form (scores ONLY): - {Consistency}
Figure 16: G-Eval prompt for NLG evaluation of the consistency of a given summarization response.
G-Eval Prompt Template for LLM Evaluation of MDS Fluency You will be given one summary written for a set of related documents. Your task is to rate the summary on one metric. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Fluency (1-3): the quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure. - 1: Poor. The summary has many errors that make it hard to understand or sound unnatural. - 2: Fair. The summary has some errors that affect the clarity or smoothness of the text, but the main points are still comprehensible. - 3: Good. The summary has few or no errors and is easy to read and follow. Evaluation Steps: 1. Read the summary. 2. Assign a score for fluency based on the Evaluation Criteria. Example: Source Text: {documents} Summary: {summary} Evaluation Form (scores ONLY): - {Fluency}
Figure 17: G-Eval prompt for NLG evaluation of the fluency of a given summarization response.
As GPT4 is expensive, we investigate use of several other commercial LLMs with the adapted G-Eval framework for evaluation on Multi-XScience and QMDSCNN, with the goal of ensuring strong alignment of resulting ratings with human evaluation. Based on computed correlations with scores from two human expert annotators on 500 samples per dataset, we find GPT-3.5-Turbo and Gemini-1.5-Flash to be most human-aligned for evaluation of model performances on Multi-XScience and QMDSCNN, respectively, and therefore use these for all reported evaluations.
Appendix D Meta-Evaluation on MDCureRM
To validate the reliability of using MDCureRM as an evaluator for MD instruction data, we conducted a meta-evaluation study to compare MDCureRM with humans. Here we provide additional details regarding the prompt and setup used in the meta-evaluation study. We utilized two annotators (graduate students in NLP working directly with LLMs) to obtain fine-grained scores from 1-5 according to the six criteria defined in MDCure on 100 randomly sampled MDCure-generated candidate instructions. Each candidate was generated using a different prompt template over distinct news article clusters from NewSHead. Annotators were informed of the purpose, aims, and intended use of the study and annotations, and informed consent was collected prior to their performing the task. No compensation was provided given the small-scale nature of the task. The meta-evaluation task was formulated using the direction shown in Figure 18, and ratings were collected via a Google form.
Meta-Evaluation Annotation Task Directions Task Description
In this task, you will evaluate the quality of instruction-answer pairs generated by a large language model (LLM). Each pair is based on a set of related news articles, with the instruction potentially being a question, command, or descriptive phrase.
Your goal is to assess how helpful each instruction-answer pair would be for instruction-tuning an LLM, specifically to improve its multi-document capabilities. To this end, you will score each instruction-answer pair across six key criteria, which focus on both the pair’s overall quality (criteria 1-3) and the pair’s effectiveness in handling multi-document content (criteria 4-6). While evaluating, prioritize the instruction over the answer, except where the criterion explicitly calls for focus on the answer (e.g., criterion #2). But loosely keep in mind the quality of the answer, especially in cases where the answer’s accuracy or depth affects the overall value of the pair. To correctly complete the task, please follow these steps: Keep this document open on the side, such that this document and the Google Form for responses are both visible at once. Read the context documents, to be aware of the information contained in each article. Note that some samples have many documents or very long documents. It is not required to scrutinize the articles in meticulous detail; rather, the annotator should obtain an understanding of the content so as to make accurate judgments regarding the instruction-answer pair. Read the proposed instruction and answer. Rate the instruction-answer pair on a scale from 1 (worst) to 5 (best) according to the following criteria: relevance, coherence & factuality, creativity, context integration, inter-document relationships, and complexity. Criteria Definitions Relevance: How well does the instruction align with the context provided by the documents? This involves assessing the pertinence of the instruction to the content and the degree to which it makes sense given the information in the documents. Coherence & Factuality: Do the instruction and answer form a coherent, logical, and factually accurate pair? This involves checking if the answer directly addresses the instruction, if the level of detail in the answer is appropriate for the instruction, and if the reasoning behind the answer is sound and well-supported by the documents. Creativity: How creative or diverse is the instruction in terms of the type of question or task it involves? This involves assessing the question type (e.g., factual, inferential, analytical, summarization, etc.) and answer format (e.g., multiple-choice, open-ended, etc.). Context Integration: How well does the instruction leverage the context provided by multiple documents? This assesses whether the instruction involves synthesizing information from various documents to form a comprehensive response. Inter-Document Relationships: How well does the instruction encourage understanding relationships between different documents? This assesses whether the instruction requires comparing, contrasting, drawing connections, or identifying discrepancies between the documents. Complexity: Does the instruction appropriately and effectively challenge the answerer’s ability to understand and leverage information across multiple documents? This evaluates how well the instruction involves higher-order thinking skills like analysis, synthesis, and evaluation.
Figure 18: G-Eval prompt for NLG evaluation of the fluency of a given summarization response.
Human Human & GPT Human & MDCureRM
Krippendorff ($\alpha$) 0.81 0.18 0.33
Table 9: IAA for human annotations, versus IAA and correlation for human and LLM annotations.
The results of the meta-evaluation study are shown in Table 9. We compute the inter-annotator agreement (IAA) among human annotators by calculating the Krippendorff’s alpha coefficient for each criterion and then taking the average over all 6 criteria. We also compute the IAA when GPT-3.5-Turbo and MDCureRM are each separately considered as an additional annotator to demonstrate the value of MDCureRM as a cost-effective, more human-aligned alternative to proprietary model scoring. Overall, we find that while GPT-3.5-Turbo and MDCureRM have weak agreement with human annotators, MDCureRM yields stronger alignment with humans versus GPT. This is consistent with demonstrated MD performance gains when using MDCureRM-filtered instruction data as opposed to GPT-filtered data for IT. We note that since MDCureRM is trained using data with target scores generated by GPT-4o, its alignment with human scoring could perhaps be improved by using human annotations of training data in future work.
Appendix E Full Experimental Results
IT Data Setting HQA WH MXS QC SEAM ZeroScrolls (ZS) Avg
ECB+ MN MSQ OA SC Avg GvRp SSFD QMsm NrQA Qspr QLTY SQLT MuSQ SpDg BkSS ZSS
Flan-T5 Base None 4.4 45.1 1.74 2.16 0.0 5.8 0.2 2.6 0.1 1.7 5.2 5.4 9.9 16.7 14.2 48.2 4.8 26.9 0.0 0.0 13.1 10.2
Unfiltered 31.7 47.5 4.04 2.35 0.0 6.7 0.1 2.7 0.2 1.9 5.3 8.6 12.4 15.5 28.7 46.6 6.2 30.0 54.3 2.4 21.0 16.1
GPT-Filtered 44.1 46.0 4.16 2.44 0.0 6.0 0.2 2.7 0.0 1.8 6.1 9.2 13.4 16.4 25.8 49.6 6.8 27.8 54.4 3.6 21.3 16.8
MDCureRM 47.3 48.3 4.22 2.58 0.0 7.5 0.3 2.7 0.0 2.1 6.6 9.2 14.5 16.1 34.6 48.2 8.1 31.2 54.4 3.4 22.6 17.9
Large None 24.4 54.6 3.19 2.78 1.1 7.9 0.1 3.1 0.3 2.5 19.5 13.5 15.5 9.4 24.4 31.0 21.1 11.8 44.4 41.1 23.2 17.3
Unfiltered 46.9 62.9 4.10 2.91 0.7 8.0 0.0 3.8 0.1 2.5 6.6 10.4 12.1 19.7 37.6 61.8 9.3 35.1 48.7 0.1 24.2 19.5
GPT-Filtered 46.3 65.1 4.13 2.88 0.5 8.3 0.0 3.9 0.1 2.6 7.2 9.9 11.9 19.8 39.1 63.4 8.3 33.8 48.8 0.0 24.2 19.7
MDCureRM 49.6 66.1 4.19 2.97 1.2 9.1 0.1 4.2 0.0 2.9 7.5 10.1 12.8 20.0 43.2 64.0 10.7 35.4 48.6 0.1 25.3 20.5
Qwen2-Ins 1.5B None 21.5 17.8 4.20 3.30 9.4 10.6 0.5 5.0 13.0 7.7 17.8 11.4 15.0 13.3 33.8 48.2 16.7 14.5 58.7 10.5 24.0 17.1
Unfiltered 32.9 30.5 4.24 3.57 15.7 12.0 0.4 5.0 16.8 10.0 14.1 9.8 15.3 15.7 40.3 48.0 15.7 16.1 53.8 10.1 23.9 18.9
GPT-Filtered 33.3 32.9 4.24 3.66 16.1 12.0 0.5 5.1 15.9 9.9 17.9 11.3 14.5 13.8 35.9 49.2 16.5 16.0 56.9 10.1 24.2 19.3
MDCureRM 37.7 34.8 4.25 3.73 16.4 12.0 0.6 5.9 18.1 10.6 19.8 12.0 14.9 13.1 38.0 55.0 16.4 19.7 54.9 12.5 25.6 20.5
7B None 30.5 39.6 4.30 3.57 5.0 11.9 0.5 6.4 13.1 7.4 21.1 12.2 15.0 13.2 33.8 72.8 19.0 29.4 16.3 6.5 23.9 18.6
Unfiltered 40.5 43.3 4.26 3.79 8.0 15.3 0.5 6.6 11.9 8.5 14.4 11.9 16.5 17.5 42.7 70.2 17.4 29.6 35.1 7.4 26.3 20.9
GPT-Filtered 42.0 44.0 4.26 3.84 8.7 15.3 0.9 6.6 12.1 8.7 11.5 12.0 16.8 18.1 43.1 67.4 15.9 28.4 54.1 19.6 28.7 22.3
MDCureRM 44.7 46.0 4.28 3.93 13.8 15.4 0.6 6.7 14.9 10.3 25.8 13.2 15.3 14.4 45.8 74.8 19.7 30.8 26.0 31.7 29.8 23.6
+PRIMERA 44.6 45.4 3.96 2.17 5.9 15.3 0.5 6.6 12.4 8.1 25.4 12.3 14.6 24.4 26.4 90.5 21.1 22.2 37.8 2.6 27.7 21.8
+QAMDen 42.6 45.8 3.92 2.37 9.5 15.3 0.5 6.4 12.4 8.8 18.8 4.5 15.7 36.0 41.5 76.2 13.6 18.4 50.7 9.5 28.5 22.3
Llama3.1-Ins 8B None 35.5 27.1 4.28 2.94 10.5 15.0 0.6 7.6 17.4 10.2 23.7 5.8 5.2 10.5 4.6 75.6 13.3 0.7 47.2 0.3 18.7 16.2
Unfiltered 37.6 34.4 3.81 4.07 15.3 16.3 0.5 7.8 18.9 11.8 20.2 14.1 17.4 21.9 52.3 50.0 19.4 25.3 32.6 32.1 28.6 22.3
GPT-Filtered 38.0 42.3 4.29 3.95 8.7 16.0 0.6 7.8 18.3 10.3 20.3 14.2 17.5 22.2 53.6 59.8 19.5 26.2 36.1 26.4 29.6 22.9
MDCureRM 44.7 43.7 4.29 4.22 16.3 16.4 0.6 7.9 18.5 11.9 19.9 14.5 17.6 22.6 52.2 68.2 19.2 27.5 34.3 33.4 30.9 24.5
70B None 53.9 38.1 4.28 3.97 25.1 21.9 0.6 6.1 11.5 13.0 21.9 14.9 18.1 25.5 45.5 66.7 23.4 42.5 61.4 43.8 36.4 27.9
Unfiltered 55.9 40.6 3.99 3.17 3.6 21.9 0.6 6.3 11.2 8.7 24.3 16.2 19.7 24.5 39.7 81.0 18.8 31.3 60.6 55.9 37.2 27.3
GPT-Filtered 57.4 41.2 3.97 3.37 4.9 22.0 0.7 6.4 10.7 8.9 24.5 15.6 17.8 21.3 41.0 81.0 18.5 30.3 69.2 54.4 37.4 27.6
MDCureRM 58.4 45.5 4.28 3.99 25.2 22.1 0.7 6.7 12.0 13.3 21.5 15.4 18.3 29.8 50.7 83.4 19.4 33.6 66.0 38.7 37.7 29.2
    Baselines PRIMERA 0.4 0.5 3.18 1.09 7.1 7.7 0.5 3.9 10.4 5.9 6.9 4.6 3.5 0.6 1.5 13.4 9.7 0.2 1.9 0.0 4.2 4.1
QAMDen 1.8 1.9 2.86 1.22 0.0 0.4 0.0 1.0 3.1 0.9 6.0 6.2 5.7 2.4 6.1 3.4 7.1 0.9 0.1 0.0 3.8 2.6
LongAlign-7B 10.4 14.3 4.15 3.75 11.5 16.5 0.0 4.1 16.8 9.8 19.5 13.5 15.5 9.4 24.4 31.0 21.1 11.8 44.4 41.1 23.2 16.5
ProLong-8B 43.6 34.3 3.86 2.46 9.5 17.4 0.4 10.7 18.0 11.2 25.6 15.2 17.6 31.6 32.2 66.2 23.2 8.6 51.8 50.5 32.2 24.3
PL+MDCure 45.8 45.1 3.93 3.40 15.3 17.3 0.5 10.5 18.9 12.5 24.6 13.9 15.9 30.7 28.2 66.7 21.6 21.7 51.4 47.1 32.2 25.4
Jamba 1.5 Mini 47.5 41.8 4.24 3.92 20.1 14.2 0.0 5.3 20.4 12.0 27.4 15.3 18.2 40.9 30.9 85.7 20.1 23.6 43.0 46.4 35.1 26.8
GPT-4o 57.5 50.0 4.50 4.25 8.9 18.6 0.3 14.2 28.9 14.1 25.7 14.8 15.9 35.3 58.8 88.6 20.7 49.2 83.5 5.6 39.8 30.8
Gemini 1.5 Pro 66.6 55.6 4.22 3.59 21.4 17.8 0.6 15.1 30.1 17.0 22.1 14.3 14.3 53.9 36.1 85.7 18.1 36.1 39.4 0.0 32.0 28.2
Table 10: Full evaluation of MDCure’d models versus baselines on 6 benchmarks in the zero-shot prompting setting. Dataset abbreviations are described in App. C.1.1.